OpenStudy (whalenjacob):

Two lines, A and B, are represented by the following equations: Line A: 2x + 2y = 8 Line B: x + y = 3 Which statement is true about the solution to the set of equations? It is (1, 2). There are infinitely many solutions. It is (2, 2). There is no solution.

OpenStudy (templeguy):

can you answer mine first?

OpenStudy (whalenjacob):

ill tryyyy

OpenStudy (templeguy):

ok

OpenStudy (frostyflame):

The solution is were these two equations intersect on a graph. So you'll want to put the equations in y = mx +b form to have y = -x + 3 and y = x +4

OpenStudy (mathstudent55):

Look carefully at both equations. Notice that the first equation has 2x and 2y, and the second equation has x and y. Divide equation A by 2. Now compare the equations again. What do you notice?